OpenStudy (anonymous):

Help. please show steps. I put this on here and got an answer but it was wrong. I know the problem is a difference of square. 256x^2-100

OpenStudy (anonymous):

a^2-b^2=(a-b)(a+b)

OpenStudy (anonymous):

256x^2-100=(16x-10)(16x+10)

OpenStudy (anonymous):

not quite done

OpenStudy (anonymous):

2(8x-5)2(8x+5)

OpenStudy (anonymous):

= 4(8x-5)(8x+5)

OpenStudy (anonymous):

4(8x-5)(8x+5)